# Hyper-optimized Quantum Lego Contraction Schedules

### v1

@@ I Introduction
The paper’s structure is as follows. In Section II we introduce stabilizer codes, quantum weight enumerators, and the quantum LEGO formalism, including our extension to the language. The details of the QL layouts are described in Section III. We report our hyper-optimization cost function, the numerical findings for contraction costs, and some details of our software tools in Section VI. Finally, we conclude with a summary of our results and our outlook in Section VII.
II Background
@@ II.1 Stabilizer codes
$|,\displaystyle\sum_{l_{j}\in\{0,1\}}V_{l_{1},\ldots,l_{n+k}}\outerproduct{l_{1},\ldots,l_{n}}{l_{n+1},\ldots,l_{n+k}}$, (3)
@@
II.2 Weight enumerator polynomials
@@
The formulas are simpler for qubit stabilizer codes, as the error basis becomes $\mathcal{P}_{n}$, and $P^{\dagger}=P$ for all Paulis. Let $\Pi_{C}=\frac{1}{2^{n-k}}\sum_{M\in S}M$ denote the projector to the code space a stabilizer code $C$ and stabilizer group $S$. Then, the Shor-Laflamme WEPs then become the stabilizer weight enumerator polynomial ($A$-type) and the normalizer weight enumerator polynomial ($B$-type):
@@
II.3 Quantum LEGO formalism
@@
II.3.1 PlanqTN implementation of quantum LEGO
@@
II.3.2 Quantum LEGO for code construction
@@
Here, we describe an efficient algorithm for calculating the final encoding tensor of stabilizer codes, utilizing symplectic parity check matrices. We follow the method described by Cao and Lackey [11, 13]. Calculating the stabilizers for the full network (or any subgraphs) is polynomial, as tracing only requires Gauss elimination and elementary row operations, which we will describe now.
@@
II.3.3 Quantum LEGO for weight enumerator polynomials
@@
II.4 Hyper-optimized tensor network contraction schedules
@@
III Quantum LEGO layouts
@@
III.1 Concatenated code, tree network
@@ III.2 Holographic code
Cao, Gullans, Lackey and Zang estimate the upper bound on contraction cost of these codes to be $O(n^{1+\alpha}),\alpha>0$, where $\alpha$ is dependent on the exact hyperbolic tesselation [17], resulting in super-polynomial speedup compared to brute force using the QL contraction.
III.3 Two-dimensional grid network
@@
III.4 Measurement state preparation and Tanner networks
@@
IV Stabilizer code tensors are sparse
@@
V Sparse stabilizer tensor cost function for stabilizer codes
Figure 7: Contraction cost function vs true contraction cost (number of polynomial multiplications counted during contractions) for (a) Cotengra’s default dense tensor cost function and (b) our sparse stabilizer tensor cost function. The dense tensor cost function, although it correlates with, fails to accurately estimate the true cost function. The core insight leveraged by our cost function is that we have a polynomial complexity algorithm to determine the number of non-zero elements in the weight enumerator polynomial tensors for stabilizer codes as described in Section V. This translates to significant speedups for most cases examined in Section VI.
As demonstrated in Section IV, the tensors in our stabilizer code tensor networks stay sparse throughout the contraction. Here, we describe a polynomial-time algorithm to determine the non-zero elements in the intermediate tensors, as well as an algorithm to determine the number of matching elements during contraction. The backbone of our algorithm is the parity check matrix calculation algorithm described in Section II.3.2.
For a given contraction tree, we can estimate the contraction cost by executing the contraction and keeping track of the PCM of intermediate tensors. For a WEP calculation, as described in Section II.3.3, at every step we contract two separate tensors on some legs. The number of operations in a contraction of two tensors is taken to be the number of polynomial multiplications. Our sparse implementation of the algorithm for WEP calculation described in Section II.3.3 for a given contraction tree is simply iterating through all the possible basis elements $e_{E_{1}\otimes E_{2}}$ in the tensor product, but only calculating the multiplication $A_{E_{1}}^{(J_{1})}(z;\Pi_{C_{1}})A_{E_{2}}^{(J_{2})}(z;\Pi_{C_{2}})$ between the two polynomials for matching elements. A matching element has the same operators on the corresponding indices defined by the contraction, or formally, for a contraction between the indices $A=\{(a_{1},a_{2})|a_{1}\in J_{1},a_{2}\in J_{2}\}$, a matching element has the property that for all $(a_{1},a_{2})\in A:(E_{1})_{a_{1}}=(E_{2})_{a_{2}}$. We count these elements using an $O(n^{3})$ algorithm below.
@@ V Sparse stabilizer tensor cost function for stabilizer codes
$|row(W\equiv\left(\begin{array}[]{c}H_{1}|_{J_{1}}\\ H_{2}|_{J_{2}}\end{array}\right))|$ $=\frac{2^{rank(H_{1}|_{J_{1}})}2^{rank(H_{2}|_{J_{2}})}}{|I|}$
@@ VII Conclusions and outlook
SST, alongside Cotengra’s hyper-optimized contraction schedules, increases the potential for QL to find new classes of codes for which the entanglement structure and optimal QL layouts are viable for WEP calculation. Our tools thus aid in code design, and we are hopeful that they will lead to the expansion of the class of practical, WEP-viable codes that have the most significant impact.
Acknowledgements.
@@ Appendix B Computing the tensor size in the Hyper-Greedy strategy
While the Hyper-Greedy strategy in our implementation uses the SST cost function for the hyper-optimizer part, it internally uses the dense tensor assumption for determining the tensor sizes Eq. 19. This could be further improved by using the rank instead, as it is used in the SST cost function described in Section V. In Fig. 11, we display the log of the true size of intermediate tensors versus the number of legs. This shows that there are plenty of cases where, using the dense assumption, the Greedy function would pick tensors with less open legs, but that might actually have a higher actual size based on the rank.
@@ Appendix C Cost function time analysis
Figure 12: Comparison of the computation time for the SST cost function vs the Cotengra default cost function for the Hyper-Greedy method.
@@
Figure 13: Comparison of the computation time for the SST cost function vs the Cotengra default cost function for the Hyper-Par method.
@@
Figure 14: The Hyper-Greedy algorithm on QL networks using the default Cotengra cost function and the sparse stabilizer tensor cost function with a 5-minute cutoff. The red dotted bars indicate the cost required to calculate the WEP using brute-force counting. The number above the bars indicates the ratio of dense tensor cost operations to SST operations, showing the speedup factor. For small examples, we compared the results to the output of the slow Optimal strategy (dashed). It is clear that Cotengra reaches close to optimal contraction trees in these cases. The bar heights are the geometric mean of the contraction costs, and the error bars are the geometric standard deviation. We choose to use the geometric mean given the skew in the data.
@@
Figure 15: The Hyper-Par algorithm on QL networks using the default Cotengra cost function and the sparse stabilizer tensor cost function with a 5-minute cutoff. The red dotted bars indicate the cost required to calculate the WEP using brute-force counting. The number above the bars indicates the ratio of dense tensor cost operations to SST operations, showing the speedup factor. For small examples, we compared the results to the output of the slow Optimal strategy (dashed). It is clear that Cotengra reaches close to optimal contraction trees in these cases. The bar heights are the geometric mean of the contraction costs, and the error bars are the geometric standard deviation. We choose to use the geometric mean given the skew in the data.
@@
Computation times for the SST and default cost functions are shown in Fig. 12 for the Hyper-Greedy method and in Fig. 13 for Hyper-Par.
@@
For Hyper-Greedy, we see that the use of the SST cost function increases computation time by at most 10 minutes across all codes, and is typically only about 1-3 minutes slower. For the smaller codes where the overall contraction cost is already low and thus the actual contraction time is a matter of seconds, the additional time from using the SST cost function may not be justified. However, for larger codes, this slight overhead can lead to large savings for the actual contraction. For example, in the $n=95$ Holographic code, using the default cost function results in an average total contraction time of 24.48 minutes (where the total contraction time includes time from the cost function to find the schedule plus the actual contraction). When using the SST cost function, even though it is slightly longer for the contraction tree optimization, the total time is only 4 minutes on average.
@@
For Hyper-Par, we see a larger range of computation times for both the SST and default cost functions. This is because the Hyper-Par strategy is a recursive ensemble of optimizers, and every level adds an extra level of uncertainty. The additional $O(n^{3})$ overhead of the SST exacerbates the effect.
@@
To evaluate the performance of the SST and default cost functions in terms of time, we ran 100 runs of each code with a 5-minute cutoff instead of a fixed number of trials. The results are shown in Fig. 14 for Hyper-Greedy and in Fig. 15 for Hyper-Par. Since Hyper-Greedy still runs quickly, we observe similar improvements to those seen in the experiment with a fixed number of trials. The SST cost function finds lower contraction costs across all codes, except for the one equal case, as expected. On the other hand, for Hyper-Par, we observe that the default cost function outperforms the SST cost function for the larger codes. This is because the SST cost function does not run as many trials as the default cost function within the 5-minute-cutoff limit.

### v2

@@ 1 Introduction
The paper’s structure is as follows. In Section 2 we introduce stabilizer codes, quantum weight enumerators, and the quantum LEGO formalism, including our extension to the language. The details of the QL layouts are described in Section 3. We report our hyper-optimization cost function, the numerical findings for contraction costs, and some details of our software tools in Section 6. Finally, we conclude with a summary of our results and our outlook in Section 7.
2 Background
@@ 2.1 Stabilizer codes
We will denote with $\Pi_{C}=\frac{1}{2^{n-k}}\sum_{M\in S}M$ the projector to the code space of a stabilizer code $C$ with stabilizer group $S$.
@@
$|,\displaystyle\sum_{l_{j}\in\{0,1\}}V_{l_{1}\ldots l_{n+k}}\outerproduct{l_{1}\ldots l_{n}}{l_{n+1}\ldots l_{n+k}}$, (3)
@@
2.2 Weight enumerator polynomials
@@
For a qubit stabilizer code $C$, with code space projector $\Pi_{C}$, the observables in Eqs. 5 and 6 become $O_{1}=O_{2}=\Pi_{C}$ in Eqs. 7 and 8 and the error basis $\mathcal{E}$ becomes $\mathcal{P}_{n}$. As $P^{\dagger}=P$ for all Paulis, the formulas become simpler, the stabilizer weight enumerator polynomial ($A$-type) and the normalizer weight enumerator polynomial ($B$-type) for qubit stabilizer code $C$ are:
@@
2.3 Quantum LEGO formalism
@@
2.3.1 PlanqTN implementation of quantum LEGO
@@
2.3.2 Quantum LEGO for code construction
@@
Here, we describe an efficient algorithm for calculating the final encoding tensor of stabilizer codes, utilizing symplectic parity check matrices.
We follow the method described by Cao and Lackey [11, 13]. Calculating the stabilizers for the full network (or any subgraphs) is polynomial, as tracing only requires Gauss elimination and elementary row operations, which we will describe now.
@@
2.3.3 Quantum LEGO for weight enumerator polynomials
@@
2.4 Hyper-optimized tensor network contraction schedules
@@
3 Quantum LEGO layouts
@@
3.1 Concatenated code, tree network
@@ 3.2 Holographic code
Cao, Gullans, Lackey and Wang estimate the upper bound on contraction cost of these codes to be $O(n^{1+\alpha}),\alpha>0$, where $\alpha$ is dependent on the exact hyperbolic tesselation [17], resulting in super-polynomial speedup compared to brute force using the QL contraction.
3.3 Two-dimensional grid network
@@
3.4 Measurement state preparation and Tanner networks
@@
4 Stabilizer code tensors are sparse
@@
5 Sparse stabilizer tensor cost function for stabilizer codes
Figure 7: Contraction cost function vs true contraction cost (number of polynomial multiplications counted during contractions) for (a) Cotengra’s default dense tensor cost function and (b) our sparse stabilizer tensor cost function. The dense tensor cost function, although it correlates with, fails to accurately estimate the true cost function. The core insight leveraged by our cost function is that we have a polynomial complexity algorithm to determine the number of non-zero elements in the weight enumerator polynomial tensors for stabilizer codes as described in Section 5. This translates to significant speedups for most cases examined in Section 6.
As demonstrated in Section 4, the tensors in our stabilizer code tensor networks stay sparse throughout the contraction. Here, we describe a polynomial-time algorithm to determine the non-zero elements in the intermediate tensors, as well as an algorithm to determine the number of matching elements during contraction. The backbone of our algorithm is the parity check matrix calculation algorithm described in Section 2.3.2.
For a given contraction tree, we can estimate the contraction cost by executing the contraction and keeping track of the PCM of intermediate tensors. For a WEP calculation, as described in Section 2.3.3, at every step we contract two separate tensors on some legs. The number of operations in a contraction of two tensors is taken to be the number of polynomial multiplications. Our sparse implementation of the algorithm for WEP calculation described in Section 2.3.3 for a given contraction tree is simply iterating through all the possible basis elements $e_{E_{1}\otimes E_{2}}$ in the tensor product, but only calculating the multiplication $A_{E_{1}}^{(J_{1})}(z;\Pi_{C_{1}})A_{E_{2}}^{(J_{2})}(z;\Pi_{C_{2}})$ between the two polynomials for matching elements. A matching element has the same operators on the corresponding indices defined by the contraction, or formally, for a contraction between the indices $A=\{(a_{1},a_{2})|a_{1}\in J_{1},a_{2}\in J_{2}\}$, a matching element has the property that for all $(a_{1},a_{2})\in A:(E_{1})_{a_{1}}=(E_{2})_{a_{2}}$. We count these elements using an $O(n^{3})$ algorithm below.
@@ 5 Sparse stabilizer tensor cost function for stabilizer codes
as the generators of the rowspace of $B$ are disjoint on the two halves of $H_{1}|J_{1}$ and $H_{2}|J_{2}$. In contrast, we define the matrix
$W\equiv\left(\begin{array}[]{c}H_{1}|_{J_{1}}\\ H_{2}|_{J_{2}}\end{array}\right),$
the rows of which will generate the union of elements generated by the two sub-PCMs. To count these, we have to remove the intersection to avoid double-counting, and hence:
$|row(W)|$ $=\frac{2^{rank(H_{1}|_{J_{1}})}2^{rank(H_{2}|_{J_{2}})}}{|I|}.$
@@ 6 Performance results
The performance results for Hyper-Greedy are shown in Fig. 8 and for Hyper-Par in Fig. 9. The average values and improvement factors are shown in Appendix D.
@@ 7 Conclusions and outlook
The SST cost function, alongside Cotengra’s hyper-optimized contraction schedules, increases the potential for QL to find new classes of codes for which the entanglement structure and optimal QL layouts are viable for WEP calculation. Our tools thus aid in code design, and we are hopeful that they will lead to the expansion of the class of practical, WEP-viable codes.
@@ Appendix B Computing the tensor size in the Hyper-Greedy strategy
While the Hyper-Greedy strategy in our implementation uses the SST cost function for the hyper-optimizer part, it internally uses the dense tensor assumption for determining the tensor sizes Eq. 18. This could be further improved by using the rank instead, as it is used in the SST cost function described in Section 5. In Fig. 11, we display the log of the true size of intermediate tensors versus the number of legs. This shows that there are plenty of cases where, using the dense assumption, the Greedy function would pick tensors with less open legs, but that might actually have a higher actual size based on the rank.
@@ Appendix C Cost function time analysis
Figure 12: Comparison of the computation time for the SST cost function vs the default dense cost function for the Hyper-Greedy method.
@@
Figure 13: Comparison of the computation time for the SST cost function vs the default dense cost function for the Hyper-Par method.
@@
Figure 14: The Hyper-Greedy algorithm on QL networks using the default dense cost function and the SST cost function with a 5-minute cutoff. The red dotted bars indicate the cost required to calculate the WEP using brute-force counting. The number above the bars indicates the ratio of dense tensor cost operations to SST operations, showing the speedup factor. For small examples, we compared the results to the output of the slow Optimal strategy (dashed). It is clear that Cotengra reaches close to optimal contraction trees in these cases. The bar heights are the geometric mean of the contraction costs, and the error bars are the geometric standard deviation. We choose to use the geometric mean given the skew in the data.
@@
Figure 15: The Hyper-Par algorithm on QL networks using the default dense cost function and the SST cost function with a 5-minute cutoff. The red dotted bars indicate the cost required to calculate the WEP using brute-force counting. The number above the bars indicates the ratio of dense tensor cost operations to SST operations, showing the speedup factor. For small examples, we compared the results to the output of the slow Optimal strategy (dashed). It is clear that Cotengra reaches close to optimal contraction trees in these cases. The bar heights are the geometric mean of the contraction costs, and the error bars are the geometric standard deviation. We choose to use the geometric mean given the skew in the data.
@@
Computation times for the SST and default dense cost functions are shown in Fig. 12 for the Hyper-Greedy method and in Fig. 13 for Hyper-Par.
@@
For Hyper-Greedy, we see that the use of the SST cost function increases computation time by at most 10 minutes across all codes, and is typically only about 1-3 minutes slower. For the smaller codes where the overall contraction cost is already low and thus the actual contraction time is a matter of seconds, the additional time from using the SST cost function may not be justified. However, for larger codes, this slight overhead can lead to large savings for the actual contraction. For example, in the $n=95$ Holographic code, using the default dense cost function results in an average total contraction time of 24.48 minutes (where the total contraction time includes time from the cost function to find the schedule plus the actual contraction). When using the SST cost function, even though it is slightly longer for the contraction tree optimization, the total time is only 4 minutes on average.
@@
For Hyper-Par, we see a larger range of computation times for both the SST and default dense cost functions. This is because the Hyper-Par strategy is a recursive ensemble of optimizers, and every level adds an extra level of uncertainty. The additional $O(n^{3})$ overhead of the SST exacerbates the effect.
@@
To evaluate the performance of the SST and default dense cost functions in terms of time, we ran 100 runs of each code with a 5-minute cutoff instead of a fixed number of trials. The results are shown in Fig. 14 for Hyper-Greedy and in Fig. 15 for Hyper-Par. Since Hyper-Greedy still runs quickly, we observe similar improvements to those seen in the experiment with a fixed number of trials. The SST cost function finds lower contraction costs across all codes, except for the one equal case, as expected. For Hyper-Par, we observe that the default dense cost function performs similarly to the SST cost function. This is because the SST cost function does not run as many trials as the default dense cost function within the 5-minute cutoff limit.
Appendix D Performance Metrics for SST and Dense Cost Functions
The average contraction costs found for the two cost functions are shown in Table 1 for Hyper-Greedy and in Table 2 for Hyper-Par. The improvement factor is also shown.
Tensor Network Qubits Avg Dense Flops Avg SST Flops Improvement Factor
Concat Rep 9 $6.20\text{\times}{10}^{01}$ $6.20\text{\times}{10}^{01}$ $1.000\text{\times}{10}^{00}$
Concat Rep 27 $3.20\text{\times}{10}^{02}$ $3.12\text{\times}{10}^{02}$ $1.026\text{\times}{10}^{00}$
Concat Rep 81 $11\text{\times}{10}^{03} 1.1$$10\text{\times}{10}^{03}$ $1.005\text{\times}{10}^{00}$
Holographic 25 $3.51\text{\times}{10}^{03}$ $2.67\text{\times}{10}^{03}$ $1.316\text{\times}{10}^{00}$
Holographic 95 $5.07\text{\times}{10}^{06}$ $2.94\text{\times}{10}^{06}$ $1.722\text{\times}{10}^{00}$
RSC 9 $2.03\text{\times}{10}^{02}$ $1.37\text{\times}{10}^{02}$ $1.474\text{\times}{10}^{00}$
RSC 25 $1.83\text{\times}{10}^{03}$ $1.26\text{\times}{10}^{03}$ $1.452\text{\times}{10}^{00}$
RSC 49 $1.37\text{\times}{10}^{04}$ $8.30\text{\times}{10}^{03}$ $1.661\text{\times}{10}^{00}$
RSC Tanner 9 $8.28\text{\times}{10}^{02}$ $7.33\text{\times}{10}^{02}$ $1.129\text{\times}{10}^{00}$
RSC Tanner 25 $2.87\text{\times}{10}^{07}$ $6.57\text{\times}{10}^{06}$ $4.374\text{\times}{10}^{00}$
RSC MSP 9 $2.21\text{\times}{10}^{03}$ $2.09\text{\times}{10}^{03}$ $1.054\text{\times}{10}^{00}$
RSC MSP 25 $5.99\text{\times}{10}^{07}$ $4.28\text{\times}{10}^{06}$ $1.399\text{\times}{10}^{01}$
Hamming Tanner 7 $1.17\text{\times}{10}^{03}$ $03\text{\times}{10}^{03} 1.1$$136\text{\times}{10}^{00}$
Hamming Tanner 15 $6.00\text{\times}{10}^{06}$ $1.86\text{\times}{10}^{06}$ $3.224\text{\times}{10}^{00}$
Hamming MSP 7 $2.54\text{\times}{10}^{03}$ $2.00\text{\times}{10}^{03}$ $1.269\text{\times}{10}^{00}$
Hamming MSP 15 $5.48\text{\times}{10}^{09}$ $1.15\text{\times}{10}^{06}$ $4.758\text{\times}{10}^{03}$
BB Tanner 18 $7.08\text{\times}{10}^{09}$ $9.65\text{\times}{10}^{08}$ $7.336\text{\times}{10}^{00}$
BB Tanner 30 $7.03\text{\times}{10}^{12}$ $6.20\text{\times}{10}^{12}$ $1.134\text{\times}{10}^{00}$
BB MSP 18 $1.23\text{\times}{10}^{13}$ $3.73\text{\times}{10}^{08}$ $3.293\text{\times}{10}^{04}$
BB MSP 30 $1.50\text{\times}{10}^{18}$ $1.08\text{\times}{10}^{13}$ $1.389\text{\times}{10}^{05}$
Table 1: Table showing the average contraction costs found for the Hyper-Greedy method by the two cost functions: dense cost function and our SST cost function. It also shows the improvement factor (dense average cost / SST average cost).
Tensor Network Qubits Avg Dense Flops Avg SST Flops Improvement Factor
Concat Rep 9 $6.20\text{\times}{10}^{01}$ $6.20\text{\times}{10}^{01}$ $1.000\text{\times}{10}^{00}$
Concat Rep 27 $3.26\text{\times}{10}^{02}$ $3.25\text{\times}{10}^{02}$ $1.004\text{\times}{10}^{00}$
Concat Rep 81 $24\text{\times}{10}^{03} 1.1$$1\text{\times}{10}^{03} 1.1$$113\text{\times}{10}^{00}$
Holographic 25 $3.40\text{\times}{10}^{03}$ $2.70\text{\times}{10}^{03}$ $1.260\text{\times}{10}^{00}$
Holographic 95 $2.58\text{\times}{10}^{06}$ $2.02\text{\times}{10}^{06}$ $1.279\text{\times}{10}^{00}$
RSC 9 $3.36\text{\times}{10}^{02}$ $1.40\text{\times}{10}^{02}$ $2.400\text{\times}{10}^{00}$
RSC 25 $1.99\text{\times}{10}^{03}$ $8.76\text{\times}{10}^{02}$ $2.274\text{\times}{10}^{00}$
RSC 49 $1.50\text{\times}{10}^{04}$ $5.07\text{\times}{10}^{03}$ $2.961\text{\times}{10}^{00}$
RSC Tanner 9 $8.18\text{\times}{10}^{02}$ $7.43\text{\times}{10}^{02}$ $1.100\text{\times}{10}^{00}$
RSC Tanner 25 $9.44\text{\times}{10}^{06}$ $7.52\text{\times}{10}^{06}$ $1.254\text{\times}{10}^{00}$
RSC MSP 9 $54\text{\times}{10}^{03} 2.2$$28\text{\times}{10}^{03}$ $1.113\text{\times}{10}^{00}$
RSC MSP 25 $2.70\text{\times}{10}^{08}$ $6.21\text{\times}{10}^{06}$ $4.347\text{\times}{10}^{01}$
Hamming Tanner 7 $27\text{\times}{10}^{03} 1.1$$15\text{\times}{10}^{03}$ $1.099\text{\times}{10}^{00}$
Hamming Tanner 15 $4.06\text{\times}{10}^{06}$ $2.54\text{\times}{10}^{06}$ $1.597\text{\times}{10}^{00}$
Hamming MSP 7 $3.22\text{\times}{10}^{03}$ $1.94\text{\times}{10}^{03}$ $1.655\text{\times}{10}^{00}$
Hamming MSP 15 $9.35\text{\times}{10}^{07}$ $3.14\text{\times}{10}^{05}$ $2.975\text{\times}{10}^{02}$
BB Tanner 18 $7.51\text{\times}{10}^{09}$ $1.05\text{\times}{10}^{08}$ $7.102\text{\times}{10}^{01}$
BB Tanner 30 $6.60\text{\times}{10}^{12}$ $91\text{\times}{10}^{11} 2.2$$265\text{\times}{10}^{01}$
BB MSP 18 $5.18\text{\times}{10}^{10}$ $3.33\text{\times}{10}^{06}$ $1.555\text{\times}{10}^{04}$
Table 2: Table showing the average contraction costs found for the Hyper-Par method by the two cost functions: dense cost function and our SST cost function. It also shows the improvement factor (dense average cost / SST average cost).